# The art of modeling nuclear reactions with weakly bound nuclei: status and perspectives

### v1

@@ 6.1 Computation of CF and ICF with CDCC
Although the CDCC method was originally envisaged as a practical tool to evaluate the elastic and breakup observables, some works have been made to use this method to obtain complete and incomplete fusion cross sections.
Some authors Hag00 ; Dia02 have proposed to identify the total fusion with the amount of flux that leaves the coupled channels set due to a short-range imaginary potential $iW_{F}(R)$, while the CF is identified with the absorption due to such potential, but restricted to bound states of the projectile only. Thus, the total fusion is computed as
$\sigma_{TF}=\frac{\pi}{2\mu E}\sum_{J_{T}}(2J_{T}+1)P_{J_{T}},$ (85)
where $E$ is the bombarding energy and $P_{J_{T}}$ is the complete fusion probability for total angular momentum $J_{T}$
@@
where $\chi^{J_{T}}_{\beta,\beta_{i}}$ are the radial solutions obtained from the coupled equations (13). In the case of CF, the expression for the cross section is identical but the sum in $\beta$ is restricted to those channels associated to bound states.
The CF obtained in this way represents a lower limit of the physical CF cross section, since one has assumed no capture of all projectile fragments (10Be and the halo neutron) from breakup channels. In reality, these events should contribute to the CF, but cannot be distinguished in this model from the capture of only one projectile fragment. In this model, the incomplete fusion $\sigma_{ICF}$ is therefore defined as the absorption from breakup channels.
@@
where $K$ is the incident wave number, $E$ the energy and $N$ is a normalization constant. Similarly, the ICF cross section is obtained from the absorption occurring in the region where one of the two fragments is inside its absorption radius while the other is outside it, i.e.,
@@
$\sigma_{\mathrm{TF}}=\frac{1}{|N|^{2}}\ \frac{K}{E}\ \ \left\langle\ \Psi^{(+)}\,\left|\,W^{(1)}+W^{(2)}\,\right|\Psi^{(+)}\ \right\rangle\,,$ (91)
@@
where $\phi_{\beta}$ and $\phi_{\gamma}$ are respectively bound and unbound states of the projectile, and $\chi_{\beta}$ and $\chi_{\gamma}$ are the corresponding wave function describing the projectile-target relative motion.
Assuming that matrix-elements of the imaginary potentials connecting bound channels to bins are negligible, Eq. (91) can be put in the form
@@
$\sigma_{\mathrm{TF}}^{B}$ $=$ $\frac{1}{|N|^{2}}\frac{K}{E}\sum_{\beta,\beta^{\prime}\,\in\,{B}}\ \left\langle\chi_{\beta}\left|W_{\beta\beta^{\prime}}^{(1)}+W_{\beta\beta^{\prime}}^{(2)}\,\right|\chi_{\beta^{\prime}}\right\rangle$ (96)
$\sigma_{TF}^{C}$ $=$ $\frac{1}{|N|^{2}}\frac{K}{E}\sum_{\gamma,\gamma^{\prime}\in\,{C}}\ \left\langle\chi_{\gamma}\left|W_{\gamma\gamma^{\prime}}^{(1)}+W_{\gamma\gamma^{\prime}}^{(2)}\,\right|\chi_{\gamma^{\prime}}\right\rangle.$ (97)
where $W_{\alpha\alpha^{\prime}}^{(i)}=\left(\phi_{\alpha}\left|{W}^{(i)}\right|\phi_{\alpha^{\prime}}\right),$ with $\alpha,\alpha^{\prime}$ standing for either $\beta,\beta^{\prime}$ or $\gamma,\gamma^{\prime}$, are the matrix-elements of the imaginary potentials.
@@
$\sigma_{TF}^{B}$ $=$ $\frac{\pi}{K^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}_{B}^{TF}(J_{T})$ (98)
$\sigma_{TF}^{C}$ $=$ $\frac{\pi}{K^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}_{C}^{TF}(J_{T}),$ (99)
@@
    $\sigma_{SCF}=\frac{\pi}{K^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}^{SCF}(J_{T}).$ (106)
@@
    $\sigma_{ICFi}=\frac{\pi}{K^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}^{ICFi}(J_{T}).$ (107)
@@
An application of this model is shown in Fig. 41, where CF (left panel) and ICF (right panel) data for the reaction 7Li+209Bi Das02 ; Das04 are compared with the predictions of the model. The upper and lower panels display the same results in logarithmic and linear scale, respectively. For the CF, the agreement with the data is very satisfactory, both above and below the barrier (indicated by the arrow). For the ICF, the separate contributions for triton capture and $\alpha$ capture are shown, with the former giving a much larger cross section. The calculations are found to reproduce well the data up to $E_{\mathrm{c.m.}}\approx 34$ MeV, but overestimate them at higher energies. A detailed discussion of these results can be found in Ref. Cor20 .
@@ 6.2 Evaluation of CF and ICF cross sections with the IAV model
As discussed in Sec. 5, the IAV model provides the total inclusive cross section corresponding to the detection of the $b$ fragment in reactions of the form $A(a,b)X$. This results from the fact that the imaginary part that appears in the expectation value of Eq. (70) accounts in principle for all processes in which the participant fragment $x$ interacts nonelastically with the target nucleus. This will include the ICF cross section, but also other NEB processes not associated with the formation of a compound nucleus of the $x+A$ system, such as target excitation. The isolation of the ICF cross section from the total NEB cross section is indeed not a non-trivial problem. An intuitive approach consists in identifying the ICF with the absorption due to a short-ranged imaginary potential. This strategy has been however barely explored so far within the IAV model.
@@
In this expression, $\sigma_{\mathrm{inel}}$ corresponds to the excitation of the projectile and/or target without dissociation (i.e., inelastic scattering). The terms $\sigma_{\mathrm{EBU}}$ and $\sigma^{(b,x)}_{\mathrm{NEB}}$ correspond to the EBU and NEB contributions. In the latter, one distinguishes the cases in which either the fragment $x$ or $b$ interacts nonelastically with the target whereas the other scatters elastically. A successful determination of the CF cross section from the decomposition (111) requires that all other quantities involved in this formula can be evaluated accurately. The pure inelastic scattering cross sections ($\sigma_{\mathrm{inel}}$) are standardly computed by means of coupled-channels calculations including low-lying collective excitations of the projectile and target. The EBU part can be accurately calculated with the continuum-discretized coupled-channels (CDCC) method. Finally, the NEB contributions can be evaluated with the IAV model, at least for projectiles with a developed two-boy structure, such as 6,7Li.
An application of the method to the 7Li+209Bi reaction is shown in the top panel of Fig. 42, adapted from Ref. Jin19b . The circles are the CF data from Ref. Das04 , the solid green line is the reaction cross section obtained from a CDCC calculation and the solid red line the calculated CF cross section inferred from Eq. (111) assuming a two-body model ($\alpha+t$) for 7Li. For comparison, a single-channel barrier penetration model (BPM) calculation (dashed lined) is shown. As can be seen, the data are largely suppressed with respect to the BPM. In contrast, the CF extracted from Eq. (111) explains very well the data. As discussed in Jin19a , the reduction with respect to the BPM is found to be mainly due to the competition due to the 209Bi(7Li, $\alpha$)X channel, which includes, among others, the $t$-ICF channel (see the bottom panel).
@@ 6.3 Application to surrogate reactions
The gradual improvement in the models oriented to the computation of ICF cross sections has driven their applicability of the so-called surrogate method (SRM). This method provides an indirect way of evaluating compound-nucleus cross sections in reactions for which the direct measurement is difficult or even not possible. An example is the extraction of neutron-induced cross sections of the form $(n,\chi)$, where $\chi$ is a given decay channel product ($\gamma$, fission fragment, etc). Following the Bohr hypothesis, it is customarily assumed that in these reactions the formation and decay of a compound nucleus take place independently of each other. To obtain information on the decay of the compound nucleus ($B^{*}$) occurring in the reaction of interest ($a+A\to B^{*}\to c+C$), one uses the alternative (surrogate) reaction $a+A\to b+B^{*}$ that involves a projectile-target combination ($a+A$) that is experimentally more accessible.
We consider as an example the extraction of $(n,\chi)$ cross sections from a $(d,pX)$ surrogate reaction. Compound nuclear reactions are properly described in the Hauser-Feshbach formalism, which considers the conservation of angular momentum $J$ and parity $\pi$. The cross section for the “desired” reaction $A(n,\chi)$ is given by
@@
where $\sigma_{J_{T},\pi}^{CN}(E_{ex},J_{T},\pi)$ is the cross section for the CN formation and $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ the branching ratio for the decay in the channel $\chi$. Note that the excitation energy $E_{ex}$ of the compound nucleus $B^{*}$ is related to the center-of-mass energy $E_{a}$ in the entrance channel through the energy needed to separate $a$ from $B$: $E_{a}=E-S_{a}(B)$. The objective of the surrogate method is to experimentally determine or constrain the decay probabilities $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$, which are often difficult to calculate accurately.
In the surrogate reaction, the same CN nucleus $B^{*}$ is formed and the decay product of interest ($\chi$) is measured in coincidence with the outgoing particle $b$. The probability for this process can be written as:
@@
where the subscript $S$ denotes the specific surrogate reaction (in this case, $A(d,p)B^{*}$), $F_{S}^{CN}(E_{ex},J_{T},\pi)$ is the probability of forming $B^{*}$ in this surrogate reaction (with specific values of $E_{ex}$, $J$ and $\pi$) and $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ is the same branching ratio appearing in the direct reaction (112). The probability $P_{S,\chi}(E_{ex})$ can be obtained experimentally as the ratio between the number of coincidences between the $b$ particle and the decay particle $\chi$, $N_{S,\chi}$, and the total number of surrogate events, $N_{S}$, i.e.:
@@
Ideally, if a reliable prediction of $F_{S}^{CN}(E_{ex},J_{T},\pi)$ is possible, with an accurate determination of $P^{exp}_{S,\chi}(E_{ex})$ for a range of energies and angles of $b$, it might be possible to extract the branching ratios $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ which can then be used to calculate the desired cross section using (112). In practice, this approach is not always feasible due to the lack of some of this required information and the approach has relied on additional approximations. In particular, most practical applications have made use of the so-called “Weisskopf-Ewing approximation” (WEA) , which assumes that the branching ratios $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ are independent of the angular momentum and spin, giving rise to the simplified cross section:
@@
So, under the validity of the WEA, the neutron $(n,\chi)$ cross section can be readily inferred from the measured probabilities for the surrogate reaction. Note also that the CN cross section must be estimated in some way, using, for example, an optical model calculation.
In practice, the validity of the WEA is not guaranteed in most cases so one needs to resort to the more general expressions (112) and (113). The probabilities $F_{S}^{CN}(E_{ex},J_{T},\pi)$ that appear in the latter can be estimated with the IAV model discussed in Sec. 5.1. This idea has been successfully applied to the 95Mo$(n,\gamma)$ reaction Rat19 . The direct $(n,\gamma$) cross section for this reaction are compared in Fig. 43 with the cross section extracted from the surrogate reaction 95Mo(d,p) using the aforementioned formalism. As can be seen, they are in excellent agreement. As also shown in this figure, the result using the WEA departs significantly from the direct measurement.
Figure 43: Application of the surrogate method to the 95Mo$(n,\gamma)$ reaction. Top (panel (A)): Calculations of the cumulative probability of forming the CN through 95Mo$(d,p)$ ($\sum F^{CN}_{{}^{95}Mo(d,p)}$). The shaded region, which spans excitation energies in the range from $E_{x}=8.55$ MeV to $E_{x}=10.65$ MeV, indicates the excitation energies over which the surrogate data are fit. The vertical dotted line represents the neutron evaporation threshold. The middle (panel (B)) histogram of the total contribution to CN formation over the shaded range in (A) as a function of angular momentum, decomposed into positive and negative parities, and normalized to one over the integration region. Bottom: The solid blue region is the $(n,\gamma)$ cross section obtained from the SRM, the red circles and black squares are the direct measurements Abr08 ; Mus76 . The uncertainty resulting from experimental data and fitting error is indicated by the shaded band. The result obtained using the Weisskopf-Ewing approximation is also shown (gold diamonds). Quoted from Ref. Rat19 , with permission from APS.

### v2

@@ 6.1 Computation of CF and ICF with CDCC
Although the CDCC method was originally envisaged as a practical tool to evaluate the elastic and breakup observables, some works have been done to use this method to obtain complete and incomplete fusion cross sections.
Some authors Hag00 ; Dia02 have proposed to identify the total fusion with the amount of flux that leaves the coupled channels set due to a short-range imaginary potential $iW_{F}(R)$, while CF is identified with the absorption due to such potential, but restricted to bound states of the projectile only. Thus, total fusion is computed as
$\sigma_{TF}=\frac{\pi}{\hbar^{2}K_{0}}\sum_{J_{T}}(2J_{T}+1)P_{J_{T}},$ (95)
where $K_{0}$ is the wavenumber of the incident channel and $P_{J_{T}}$ is the complete fusion probability for total angular momentum $J_{T}$
@@
where $\chi^{J_{T}}_{\beta,\beta_{i}}$ are the radial solutions obtained from the coupled equations (14). In the case of CF, the expression for the cross section is identical but the sum in $\beta$ is restricted to those channels associated with bound states.
The CF obtained in this way represents a lower limit of the physical CF cross section, since one has assumed no capture of all projectile fragments from breakup channels. In reality, these events should contribute to the CF, but cannot be distinguished in this model from the capture of only one projectile fragment. In this model, the incomplete fusion $\sigma_{ICF}$ is therefore defined as the absorption from breakup channels.
@@
where $K_{0}$ is the incident wave number, and $E$ is the energy. Similarly, the ICF cross section is obtained from the absorption occurring in the region where one of the two fragments is inside its absorption radius while the other is outside it, i.e.,
@@
$\sigma_{\mathrm{TF}}=\frac{2\mu}{\hbar^{2}K_{0}}\left\langle\ \Psi^{(+)}\,\left|\,W^{(1)}+W^{(2)}\,\right|\Psi^{(+)}\ \right\rangle\,,$ (101)
@@
where $\phi_{\beta}$ and $\phi_{\gamma}$ are, respectively, the bound and unbound states of the projectile, and $\chi_{\beta}$ and $\chi_{\gamma}$ are the corresponding wave function describing the projectile-target relative motion.
Assuming that matrix elements of the imaginary potentials connecting bound channels to bins are negligible, Eq. (101) can be put in the form
@@
$\sigma_{\mathrm{TF}}^{B}$ $=$ $\frac{2\mu}{\hbar^{2}K_{0}}\sum_{\beta,\beta^{\prime}\,\in\,{B}}\ \left\langle\chi_{\beta}\left|W_{\beta\beta^{\prime}}^{(1)}+W_{\beta\beta^{\prime}}^{(2)}\,\right|\chi_{\beta^{\prime}}\right\rangle$ (106)
$\sigma_{TF}^{C}$ $=$ $\frac{2\mu}{\hbar^{2}K_{0}}\sum_{\gamma,\gamma^{\prime}\in\,{C}}\ \left\langle\chi_{\gamma}\left|W_{\gamma\gamma^{\prime}}^{(1)}+W_{\gamma\gamma^{\prime}}^{(2)}\,\right|\chi_{\gamma^{\prime}}\right\rangle.$ (107)
where $W_{\alpha\alpha^{\prime}}^{(i)}=\left(\phi_{\alpha}\left|{W}^{(i)}\right|\phi_{\alpha^{\prime}}\right),$ with $\alpha,\alpha^{\prime}$ standing for either $\beta,\beta^{\prime}$ or $\gamma,\gamma^{\prime}$, are the matrix elements of the imaginary potentials.
@@
$\sigma_{TF}^{B}$ $=$ $\frac{\pi}{K_{0}^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}_{B}^{TF}(J_{T})$ (108)
$\sigma_{TF}^{C}$ $=$ $\frac{\pi}{K_{0}^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}_{C}^{TF}(J_{T}),$ (109)
@@
    $\sigma_{SCF}=\frac{\pi}{K_{0}^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}^{SCF}(J_{T}).$ (116)
@@
    $\sigma_{ICFi}=\frac{\pi}{K_{0}^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}^{ICFi}(J_{T}).$ (117)
@@
An application of this model is shown in Fig. 45, where CF (left panel) and ICF (right panel) data for the reaction 7Li+209Bi Das02 ; Das04 are compared with the predictions of the model. The upper and lower panels display the same results in logarithmic and linear scale, respectively. For CF, the agreement with the data is very satisfactory, both above and below the barrier (indicated by the arrow). For ICF, the separate contributions for triton capture and $\alpha$ capture are shown, with the former giving a much larger cross section. The calculations are found to reproduce well the data up to $E_{\mathrm{c.m.}}\approx 34$ MeV, but overestimate them at higher energies. A detailed discussion of these results can be found in Ref. Cor20 .
@@ 6.2 Evaluation of CF and ICF cross sections with the IAV model
As discussed in Sec. 5, the IAV model provides the total inclusive cross section corresponding to the detection of the $b$ fragment in reactions of the form $A(a,b)X$. This results from the fact that the imaginary part that appears in the expectation value of Eq. (78) accounts in principle for all processes in which the participant fragment $x$ interacts nonelastically with the target nucleus. This will include the ICF cross section, but also other NEB processes not associated with the formation of a compound nucleus of the $x+A$ system, such as target excitation. The isolation of the ICF cross section from the total NEB cross section is indeed not a trivial problem. An intuitive approach consists in identifying the ICF with the absorption due to a short-ranged imaginary potential.
@@
In this expression, $\sigma_{\mathrm{inel}}$ corresponds to the excitation of the projectile and/or target without dissociation (i.e., inelastic scattering). The terms $\sigma_{\mathrm{EBU}}$ and $\sigma^{(b,x)}_{\mathrm{NEB}}$ correspond to the elastic breakup (EBU) and nonelastic breakup (NEB) contributions already discussed in Sec. 5. In the latter, one distinguishes the cases in which either the fragment $x$ or $b$ interacts nonelastically with the target whereas the other scatters elastically. A successful determination of the CF cross section from the decomposition (121) requires that all other quantities involved in this formula can be evaluated accurately. The pure inelastic scattering cross sections ($\sigma_{\mathrm{inel}}$) are standardly computed by means of coupled-channels calculations including low-lying collective excitations of the projectile and target. The EBU part can be accurately calculated with the continuum-discretized coupled-channels (CDCC) method. Finally, the NEB contributions can be evaluated with the IAV model, at least for projectiles with a developed two-boy structure, such as 6,7Li.
An application of the method to the 7Li+209Bi reaction is shown in the top panel of Fig. 46, adapted from Ref. Jin19b . The circles are the CF data from Ref. Das04 , the solid green line is the reaction cross section obtained from a CDCC calculation and the solid red line is the calculated CF cross section inferred from Eq. (121) assuming a two-body model ($\alpha+t$) for 7Li. For comparison, a single-channel barrier penetration model (BPM) calculation (dashed lined) is shown. As can be seen, the data are largely suppressed with respect to the BPM. In contrast, the CF extracted from Eq. (121) explains very well the data. As discussed in Jin19a , the reduction with respect to the BPM is found to be mainly due to the competition due to the 209Bi(7Li, $\alpha$)X channel, which includes, among others, the $t$-ICF channel (see the bottom panel).
@@ 6.3 Application to surrogate reactions
The gradual improvement in models oriented to the computation of ICF cross sections has driven their applicability to the so-called surrogate method (SRM). This method provides an indirect way of evaluating compound-nucleus cross sections in reactions for which the direct measurement is difficult or even unpossible. An example is the extraction of neutron-induced cross sections of the form $(n,\chi)$, where $\chi$ is a given decay channel product ($\gamma$, fission fragment, etc). Following the Bohr hypothesis, it is customarily assumed that in these reactions the formation and decay of a compound nucleus take place independently of each other. To obtain information on the decay of the compound nucleus ($B^{*}$) that occurs in the reaction of interest ($n+A\to B^{*}\to\chi+C$), one uses the alternative (surrogate) reaction $d+D\to b+B^{*}\to b+\chi+C$ that involves a projectile-target combination ($d+D$) that is experimentally more accessible. For example, one can use the stripping reaction $d+A\to p+B^{*}\to p+\chi+C$, in which $p$ and $\chi$ are measured in coincidence.
Compound nuclear reactions are adequately described in the Hauser-Feshbach formalism, which considers the conservation of angular momentum $J$ and parity $\pi$. The cross section for the “desired” reaction $A(n,\chi)C$ is given by
@@
where $\sigma_{J_{T},\pi}^{CN}(E_{ex},J_{T},\pi)$ is the cross section for the CN formation and $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ the branching ratio for the decay to channel $\chi$. The objective of the surrogate method is to experimentally determine the decay probabilities $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$, which are often difficult to calculate accurately.
In the surrogate reaction, $d+D\to b+B^{*}\to b+\chi+C$, the same CN nucleus $B^{*}$ is formed and the decay product of interest ($\chi$) is measured in coincidence with the outgoing particle $b$. The probability for this process can be written as:
@@
where the subscript $S$ denotes the specific surrogate reaction, $F_{S}^{CN}(E_{ex},J_{T},\pi)$ is the probability of forming $B^{*}$ in this surrogate reaction (with specific values of $E_{ex}$, $J$ and $\pi$) and where $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ are the same branching ratios appearing in equation (122). The probability $P_{S,\chi}(E_{ex})$ can be obtained experimentally as the ratio between the number of coincidences between the $b$ particle and the decay particle $\chi$, $N_{S,\chi}$, and the total number of surrogate events, $N_{S}$, i.e.:
@@
Ideally, if a reliable prediction of $F_{S}^{CN}(E_{ex},J_{T},\pi)$ is possible, with an accurate determination of $P^{exp}_{S,\chi}(E_{ex})$ for a range of energies and angles of $b$, it might be possible to extract the branching ratios $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ which can then be used to calculate the desired cross section using Eq. (122). In practice, this approach is not always feasible due to the lack of some of this required information and the approach has relied on additional approximations. In particular, early applications made use of the so-called “Weisskopf-Ewing approximation” , which assumes that the branching ratios $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ are independent of the angular momentum and spin, giving rise to the simplified cross section:
@@
So, under the validity of theWeisskopf-Ewing approximation, the neutron $(n,\chi)$ cross section can be readily inferred from the measured probabilities for the surrogate reaction. Note also that the CN cross section must be estimated in some way, using, for example, an optical model calculation.
In practice, the Weisskopf-Ewing approximation is rarely justified in most cases so one needs to resort to the more general expressions (122) and (123). The probabilities $F_{S}^{CN}(E_{ex},J_{T},\pi)$ that appear in the latter can be estimated with the IAV model discussed in Sec. 5.1. This idea has been successfully applied to the 95Mo$(n,\gamma)$ reaction Rat19 . The direct $(n,\gamma$) cross section for this reaction are compared in Fig. 47 with the cross section extracted from the surrogate reaction 95Mo(d,p) using the aforementioned formalism. As can be seen, they are in excellent agreement. As also shown in this figure, the result using the Weisskopf-Ewing approximation departs significantly from the direct measurement.
Figure 47: Application of the surrogate method to the 95Mo$(n,\gamma)$ reaction. Top (panel (A)): Calculations of the cumulative probability of forming the CN through 95Mo$(d,p)$ ($\sum F^{CN}_{{}^{95}Mo(d,p)}$). The shaded region, which spans excitation energies in the range from $E_{x}=8.55$ MeV to $E_{x}=10.65$ MeV, indicates the excitation energies over which the surrogate data are fit. The vertical dotted line represents the neutron evaporation threshold. Middle (panel (B)): histogram of the total contribution to CN formation over the shaded range in (A) as a function of angular momentum, decomposed into positive and negative parities, and normalized to one over the integration region. Bottom: The solid blue region is the $(n,\gamma)$ cross section obtained from the SRM, the red circles and black squares are the direct measurements Abr08 ; Mus76 . The uncertainty resulting from experimental data and fitting error is indicated by the shaded band. The result obtained using the Weisskopf-Ewing approximation is also shown (gold diamonds). Quoted from Ref. Rat19 , with permission from APS.